OpenStudy (anonymous):

solve for b: squareroot(4b/3)=16

OpenStudy (anonymous):

square on both sides , so u hav, 4b/3=256 b=192

OpenStudy (anonymous):

isnt it b=64

OpenStudy (anonymous):

*3 4b/3=256*3 b=256*3/4=192 your right